# Neutron-nucleus dynamics simulations for quantum computers

### 22 Feb 2024 v1

@@ III.0.2 Binary encoding
$|0\rangle$to $|000\rangle$,
$|1\rangle$to $|001\rangle$,
$|2\rangle$to $|010\rangle$,
$|3\rangle$to $|011\rangle$,
$|4\rangle$to $|100\rangle$,
$|5\rangle$to $|101\rangle$,
$|6\rangle$to $|110\rangle$,
$|7\rangle$to $|111\rangle$.
@@
$P^{0}$ $\coloneqq|0\rangle\!\langle 0|=\frac{1}{2}(I+Z),$
$P^{1}$ $\coloneqq|1\rangle\!\langle 1|=\frac{1}{2}(I-Z).$ (39)
@@
Note that the summation over $i$ terminates if $m+k>N-1$.
@@ III.0.3 Gray encoding
$|0\rangle$to $|000\rangle$,
$|1\rangle$to $|100\rangle$,
$|2\rangle$to $|110\rangle$,
$|3\rangle$to $|010\rangle$,
$|4\rangle$to $|011\rangle$,
$|5\rangle$to $|111\rangle$,
$|6\rangle$to $|101\rangle$,
$|7\rangle$to $|001\rangle$.
@@
Note that the summation over $k$ terminates if $m+k>N-1$.
@@ IV.1 Variational principle
For completeness, we summarize the variational principle that underpins the Variational Quantum Eigensolver (VQE). Given a Hamiltonian, the minimum energy eigenstate is called the ground state $|\psi_{g}\rangle$ and its energy $E_{g}$ is called the ground-state energy222 We note that for a model space restricted to a given total angular momentum $J$ (spin) and parity $\pi$ of the nucleus, the minimum energy eigenstate provides the lowest state for the given $J^{\pi}$ and coincides with the ground state only if the ground state has the same spin-parity. For example, in the present study of Carbon isotopes, the lowest ${\frac{1}{2}}^{+}$ state is the ground state only for ${}^{15}$ (the composite system for n+${}^{14}$) and ${}^{19}$ (the composite system for n+${}^{18}$), while the other Carbon isotopes under consideration have a ground state of different spin-parity. . More precisely, given a Hamiltonian $H$, the ground-state energy and the ground state are defined as:
$E_{g}$ $\coloneqq\min\limits_{|\psi\rangle}\ \langle\psi|H|\psi\rangle,$
$|\psi_{g}\rangle$ $\coloneqq\underset{|\psi\rangle}{\operatorname{argmin}}\ \langle\psi|H|\psi\rangle.$ (59)
@@ IV.2.1 One-hot ansatz
Thus, for a truncation parameter $N$, the state is parameterized by $N-1$ parameters. The encoded state can be generated recursively using $R_{y}$ rotation gates and $\operatorname{CNOT}$ gates, as seen in Fig. 2.
@@ V Encoding techniques and trade-offs
We note that, in the present study, the Hamiltonian for $K=0$ (diagonal potential) is tridiagonal due to the kinetic energy term. As a result, the entries for $K=1$ are used for $K=0$. On the other hand, the most general results for a Hamiltonian matrix of $12K+$ 1 bandwidth (that permits a diagonal matrix) are summarized in Table 1 for the one-hot encoding and in Table 2 for the binary and Gray encodings.
@@ V.1 Number of Pauli terms
As seen in Eq. (29), the encoded Hamiltonian for truncation parameters $N,K$ is given by (for simplicity, we take $K\geq 1$; the case for $K=0$ can be dealt with separately):
@@
with the $k$ summation terminating when $m+k>N-1$. This results from the truncation of the matrix to size $N\times N$.
In the first summation, the identity operator ($I^{\otimes N}$) and the $N$ individual $Z$ operators give a total of $N+1$ Pauli terms (e.g., for three qubits, the terms are $III$, $ZII$, $IZI$, and $IIZ$). For the second term, we split the $m$ summation into two parts: one with $m<N-K$ and $m\geq N-K$. In the first part, each $k$ summation goes from $1$ to $K$, contributing $2K$ terms. Thus, this part adds a total of $2K(N-K)$ terms. In the second part, the $k$ summation is truncated before it reaches $K$. This contributes a total of $2[(K-1)+(K-2)+\ldots+1]$. Thus, the total of the number of Pauli terms is
@@
For the example of $N=4$ and $K=2$, we see that the number of terms is $15$, confirmed by (30). The case where $K=0$ is the same case as when $K=1$ since the kinetic energy terms already account for two off-diagonal terms. Note that each of these terms acts on $N$ qubits. We note that for $K>N/2$, we see that the number of Pauli terms is $\mathcal{O}(N^{2})$.
$H_{N,K}$
One-hot
@@
Qubits
$N$
Pauli Terms
$1+N+2NK-K(K+1)$
QC Sets
$3$
Ansatz
$2$ one-qubit gates + $(2N-3)$ two-qubit gates
Table 1: Number of qubits, Pauli terms, and qubit-wise commuting sets for the one-hot code, for a general Hamiltonian matrix of $12K+$ 1 bandwidth. In the present study, the Hamiltonian for $K=0$ (diagonal potential) is tridiagonal because of the kinetic energy term, in which case the entries for $K=1$ should be used. More details can be found in Sec. VI.
V.1.1 Binary and Gray encoding
@@
where $L^{i}_{j}=B^{i}_{j}$ for the binary encoding and $L^{i}_{j}=G^{i}_{j}$ for the Gray encoding. For simplicity, we take $K\geq 1$; the case for $K=0$ can be dealt with separately.
@@
$|H(N,K)|=\begin{cases}2^{n}+n2^{n-1}&K=0\\ 2^{n}+2^{n-1}\sum\limits_{k=1}^{K}\overline{n}_{k}&1\leq K\leq 2^{n-1}\\ 2^{n-1}(1+2^{n})&K>2^{n-1},\end{cases}$ (66)
as depicted in Fig. 4 and where $\overline{n}_{k}\coloneqq n-\lceil\operatorname{log}_{2}(k)\rceil$. The proof for this can be found in Lemma A.10. For the example of $N=4,K=2$, we see that the number of terms is $10$, confirmed by Eqs. (45) and (58). For $K>N/2$, we see that the number of Pauli terms is $\mathcal{O}(N^{2})$.
@@
$H_{N,K}$
Binary Gray
$K\leq N/2$
$K>N/2$
$K\leq N/2$
$K>N/2$
Qubits
$\operatorname{log}_{2}(N)$
$\operatorname{log}_{2}(N)$
Pauli Terms
$2^{n}+2^{n-1}\sum\limits_{k=1}^{K}\overline{n}_{k}$
$2^{n-1}(1+2^{n})$
$2^{n}+2^{n-1}\sum\limits_{k=1}^{K}\overline{n}_{k}$
$2^{n-1}(1+2^{n})$
QC Sets
$1+\sum\limits_{k=1}^{K}2^{|b(\bar{k})|}[1-2^{-\overline{n}_{k}}]$
$\frac{1}{2}\left(1+3^{n}\right)$
$1+\sum\limits_{k=1}^{K}\overline{n}_{k}2^{|g(k-1)|}$
$\frac{1}{2}\left(1+3^{n}\right)$
DGC Sets
$1+\sum\limits_{k=1}^{K}\overline{n}_{k}$
$2^{n}$
$1+\sum\limits_{k=1}^{K}\overline{n}_{k}$
$2^{n}$
2QG in Diag. Unitary
$\frac{1}{2}\sum\limits_{k=1}^{K}\overline{n}_{k}\left[2|b(\overline{k})|-1-\overline{n}_{k}\right]$
$1+2^{n-1}(n-2)$
$\sum\limits_{k=1}^{K}\overline{n}_{k}|g_{k-1}|$
$1+2^{n-1}(n-2)$
Ansatz
$nL$ one-qubit gates + $(n-1)L$ two-qubit gates
$nL$ one-qubit gates + $(n-1)L$ two-qubit gates
Table 2: Number of qubit, Pauli terms, qubit-wise commuting sets and distance-grouped commuting sets for both the binary and Gray code, for a general Hamiltonian matrix of $12K+$ 1 bandwidth. Here we use $N=2^{n}$ and $\overline{n}_{k}\coloneqq n-\lceil\operatorname{log}_{2}(k)\rceil$. For the ansatz, $L$ refers to the number of layers, which is chosen large enough beforehand. In the present study, the Hamiltonian for $K=0$ (diagonal potential) is tridiagonal because of the kinetic energy term, in which case the entries for $K=1$ should be used. Notably, for $K>N/2$, there is a saturation in all quantities, implying that larger diagonal width (better approximation) can be handled without an increase in the complexity of the problem. More details can be found in Sec. VI.
@@ V.2 Number of commuting sets
In addition, in this paper, we introduce a new strategy that is based on grouping Pauli terms in terms of the distance operators defined in Appendix A. We refer to this grouping as distance-grouped commutativity (DGC). While the precise structure is not relevant here, it leads to a more optimal set of Pauli operators as compared to the qubit-wise scheme. However, the diagonalizing unitary, while simple conceptually, is no longer a tensor-product of individual Pauli operators (see Appendix A for further details).
@@ V.2.4 Binary/Gray encoding: Distance-grouped commutativity
=$=\begin{cases}0.5n(n-1)&K=0\\ 0.5\sum\limits_{k=1}^{K}\overline{n}_{k}\left[2|b(\overline{k})|-1-\overline{n}_{k}\right]&1\leq K\leq 2^{n-1}\\ 1+2^{n-1}(n-2)&K>2^{n-1},\end{cases}$ (80)
@@ VI Quantum simulations and discussions
In this section, we perform quantum simulations using the Gray and one-hot encodings, with and without noise effects, and we discuss the results. The aim is to illustrate if nuclear problems with band-diagonal Hamiltonian matrices can be solved on current quantum devices and to compare the two types of encodings. Specifically, we study the lowest $\frac{1}{2}^{+}$ bound state in the n+${}^{10}$, n+${}^{12}$, and n+${}^{14}$, using an exponential potential given in Eq. (13). We also perform quantum simulations for the lowest $\frac{1}{2}^{+}$ orbit in the ab initio deduced n-$\alpha$ local potential [^3] We note that an optical potential, such as the one derived in the Green’s function approach [23], provides a mean field that, for n+$\alpha$, yields a negative-energy $\frac{1}{2}^{+}$ orbit, occupied by the protons and neutron of the $\alpha$ particle, and is associated with the physics of a hole in ${}^{4}$e, that is, with the ground state of ${}^{3}$e. The next $\frac{1}{2}^{+}$ eigensolution (in increasing energy) corresponds to a scattering state in ${}^{5}$e..
Figure 9: Energy of the lowest $\frac{1}{2}^{+}$ state in (a) n+${}^{10}$, (b) n+${}^{12}$, (c) n+${}^{14}$, (d) n+${}^{16}$, and (e) n+${}^{18}$, vs. the model-space size $N$ for $V_{\rm E}$ (13) (blue filled) and its $K=3$ approximation (purple open, for $N$ qubits; red open, for $n=\log_{2}N$ qubits), as compared to experiment (dashed line) [43, 44, 45]. Shaded area provides the error associated with the $K=3$ approximation (the small error band for n+${}^{10}$ is not visible on the plot).
@@ VI.1 Exponential potentials for neutron-Carbon dynamics
$A$ $E_{\rm expt}$ [MeV] $E_{\rm th}$ [MeV] $E_{\rm th}^{K=3}$ [MeV] $\frac{1}{\hbar\omega}V_{0}$ $c^{-\frac{1}{2}}$
n+${}^{10}$ $10$ $-6.78$ $-6.78$ $-6.74$ $-0.650$ $5.43$
n+${}^{12}$ $12$ $-1.86$ $-1.86$ $-1.74$ $-0.283$ $5.35$
n+${}^{14}$ $14$ $-1.22$ $-1.22$ $-1.04$ $-0.242$ $5.0$
n+${}^{16}$ $16$ $-0.52$ $-0.52$ $-0.23$ $-0.175$ $4.7$
n+${}^{18}$ $18$ $-0.58$ $-0.59$ $-0.30$ $-0.192$ $4.6$
@@
For Gaussian-like exponential potentials, the truncation parameter $K$ needs to be odd, since for even $K$ values, the potential curves downward with increasing distance and leads to spurious bound states. Furthermore, the $K=3$ case provides a reasonable approximation, as illustrated by the $K=3$ error band in Fig. 9. This error is smaller for the lighter isotopes, where the $\frac{1}{2}^{+}$ state is more deeply bound (larger binding energy) compared to the one for n+${}^{16}$ and n+${}^{18}$ (see also $E_{\rm th}^{K=3}$ in Table 3). It is important to note that weakly bound states converge at a slower rate compared to states with larger binding energies. As shown in Fig. 9, the $N=8$ energy (or 3 qubits for the Gray encoding) closely agrees with the exact value for n+${}^{10}$, but larger model spaces are needed for the other systems. The model-space size requirements become even larger for resonances. This suggests that the use of the Gray encoding that can reach large model spaces with fewer qubits becomes advantageous for weakly bound states and resonances. To illustrate the advantages of the Gray encoding compared to the one-hot encoding, we perform quantum simulations for n+${}^{14}$ using both encodings, which is detailed below.
VI.2 Ab initio deduced local optical potential for n$+\alpha$
@@
Figure 10: (a) Ab initio nonlocal n+${}^{4}$e potential as a function of the distance $r$ from the center-of-mass of the $\alpha$ particle for the $S_{\frac{1}{2}}$ partial wave, calculated in the SA-NCSM/GF with $\hbar\omega=16$ MeV and 15 HO shells, for zero projectile energy (see also Fig. 4 in Ref. [23] for other energies and partial waves). (b) & (c) Approximate potentials for different truncation parameters $K$, as compared to the ab initio deduced local potential labeled as “Local $V$” (black, solid) for (b) $\hbar\omega=12$ MeV and (c) $\hbar\omega=16$ MeV. The corresponding exact lowest eigenvalue $E_{0}$ is also shown (gray, dotted).
@@
with the corresponding lowest eigenvalue of the original nonlocal potential $E_{0}=-20.84$ MeV. Hence, across the range of $\hbar\omega=12$-16 MeV, the ab initio potential yields the lowest $1/2^{+}$ orbit at energy $E_{0}=-19.8\pm 1.0$ MeV, which closely agrees with the corresponding ${}^{3}$e experimental energy of $-20.58$ MeV associated with a neutron removal from the ${}^{4}$e target.
Figure 11: Quantum simulation using the Gray encoding for the energy of the lowest $\frac{1}{2}^{+}$ state for n$+^{10}$ modeled by the $V_{\rm E}$ exponential potential Eq. (13) with (a) $N=8$ ($n=3$ qubits) and (b) $N=16$ ($n=4$ qubits), with $K=3$ and for different types of simulations detailed in Sec. VI.3, as compared to the theoretical energy $E_{\rm th}^{K=3}$ labeled as “True Value” (for the Hamiltonian parameters and $E_{\rm th}^{K=3}$, see Table 3). The inset plots are the last 200 iterations.
Figure 12: Quantum simulation using the Gray encoding for the energy of the lowest $\frac{1}{2}^{+}$ state for n$+^{12}$ modeled by the $V_{\rm E}$ exponential potential Eq. (13) with (a) $N=8$ ($n=3$ qubits) and (b) $N=16$ ($n=4$ qubits), with $K=3$ and for different types of simulations detailed in Sec. VI.3, as compared to the theoretical energy $E_{\rm th}^{K=3}$ labeled as “True Value” (for the Hamiltonian parameters and $E_{\rm th}^{K=3}$, see Table 3). The inset plots are the last 200 iterations.
@@
Figure 13: Quantum simulation using the Gray encoding for the energy of the lowest $\frac{1}{2}^{+}$ state for n$+^{14}$ modeled by the $V_{\rm E}$ exponential potential Eq. (13) with (a) $N=8$ ($n=3$ qubits) and (b) $N=16$ ($n=4$ qubits), with $K=3$ and for different types of simulations detailed in Sec. VI.3, as compared to the theoretical energy $E_{\rm th}^{K=3}$ labeled as “True Value” (for the Hamiltonian parameters and $E_{\rm th}^{K=3}$, see Table 3). The inset plots are the last 200 iterations.

### 15 Mar 2026 v3

@@ III.0.2 Binary encoding
$|0\rangle$ $\to$ $|000\rangle$,
$|1\rangle$ $\to$ $|001\rangle$,
$|2\rangle$ $\to$ $|010\rangle$,
$|3\rangle$ $\to$ $|011\rangle$,
$|4\rangle$ $\to$ $|100\rangle$,
$|5\rangle$ $\to$ $|101\rangle$,
$|6\rangle$ $\to$ $|110\rangle$,
$|7\rangle$ $\to$ $|111\rangle$.
@@
$P^{0}$ $\coloneqq|0\rangle\!\langle 0|=\frac{1}{2}(I+Z),$ (39)
$P^{1}$ $\coloneqq|1\rangle\!\langle 1|=\frac{1}{2}(I-Z).$
@@
Note that the summation over $i$ terminates if m+k>N1m+k>N-1.
@@ III.0.3 Gray encoding
$|0\rangle$ $\to$ $|000\rangle$,
$|1\rangle$ $\to$ $|100\rangle$,
$|2\rangle$ $\to$ $|110\rangle$,
$|3\rangle$ $\to$ $|010\rangle$,
$|4\rangle$ $\to$ $|011\rangle$,
$|5\rangle$ $\to$ $|111\rangle$,
$|6\rangle$ $\to$ $|101\rangle$,
$|7\rangle$ $\to$ $|001\rangle$.
@@
Note that the summation over $k$ terminates if m+k>N1m+k>N-1.
@@ IV.1 Variational principle
For completeness, we summarize the variational principle that underpins the Variational Quantum Eigensolver (VQE). Given a Hamiltonian, the minimum energy eigenstate is called the ground state $|\psi_{g}\rangle$ and its energy $E_{g}$ is called the ground-state energy [^2] We note that for a model space restricted to a given total angular momentum $J$ (spin) and parity $\pi$ of the nucleus, the minimum energy eigenstate provides the lowest state for the given $J^{\pi}$ and coincides with the ground state only if the ground state has the same spin-parity. For example, in the present study of Carbon isotopes, the lowest $\frac{1}{2}$$^{+}$ state is the ground state only for 15C (the composite system for n+14C) and 19C (the composite system for n+18C), while the other Carbon isotopes under consideration have a ground state of different spin-parity. . More precisely, given a Hamiltonian $H$, the ground-state energy and the ground state are defined as:
[…]
$|\psi_{g}\rangle$ $\coloneqq\underset{|\psi\rangle}{\operatorname{argmin}}\ \langle\psi|H|\psi\rangle.$
@@ IV.2.1 One-hot ansatz
Thus, for a truncation parameter $N$, the state is parameterized by N1N-1 parameters. The encoded state can be generated recursively using $R_{y}$ rotation gates and $\operatorname{CNOT}$ gates, as seen in Fig. 2.
@@ V Encoding techniques and trade-offs
We note that, in the present study, the Hamiltonian for $K=0$ (diagonal potential) is tridiagonal due to the kinetic energy term. As a result, the entries for $K=1$ are used for $K=0$. On the other hand, the most general results for a Hamiltonian matrix of $2K+1$ bandwidth (that permits a diagonal matrix) are summarized in Table 1 for the one-hot encoding and in Table 2 for the binary and Gray encodings.
@@ V.1 Number of Pauli terms
V.1.1 One-hot encoding
As seen in Eq. (29), the encoded Hamiltonian for truncation parameters $N,K$ is given by
@@ V.1.1 One-hot encoding
with the $k$ summation terminating when m+k>N1m+k>N-1. The above equation is true for $K\geq 1$; the case for $K=0$ is the same as $K=1$, since the kinetic energy term is tridiagonal. This results from the truncation of the matrix to size $N\times N$.
In the first summation, the identity operator ($I^{\otimes N}$) and the $N$ individual $Z$ operators give a total of $N+1$ Pauli terms (e.g., for three qubits, the terms are $III$, ZIIZII, IZIIZI, and $IIZ$). For the second term, we split the $m$ summation into two parts: one with m<NKm<N-K and $m\geq N-K$. In the first part, each $k$ summation goes from $1$ to $K$, contributing $2K$ terms. Thus, this part adds a total of 2K(NK)2K(N-K) terms. In the second part, the $k$ summation is truncated before it reaches $K$. This contributes a total of $2[(K-1)+(K-2)+\ldots+1]$. Thus, the total of the number of Pauli terms is
@@
For the example of $N=4$ and $K=2$, we see that the number of terms is $15$, confirmed by (30). Note that each of these terms acts on $N$ qubits. We note that for $K>N/2$, we see that the number of Pauli terms is $\mathcal{O}(N^{2})$.
$H_{N,K}$ One-hot
@@
Qubits NN
Pauli Terms 1+N+2NKK(K+1)1+N+2NK-K(K+1)
QC Sets 33
Ansatz 22 one-qubit gates + (2N3)(2N-3) two-qubit gates
Table 1: Number of qubits, Pauli terms, and qubit-wise commuting sets for the one-hot code, for a general Hamiltonian matrix of $2K+1$ bandwidth. In the present study, the Hamiltonian for $K=0$ (diagonal potential) is tridiagonal because of the kinetic energy term, in which case the entries for $K=1$ should be used. More details can be found in Sec. VI.
V.1.2 Binary and Gray encoding
@@
where $L^{i}_{j}=B^{i}_{j}$ for the binary encoding and $L^{i}_{j}=G^{i}_{j}$ for the Gray encoding. The above equation is true for $K\geq 1$; the case for $K=0$ is the same as $K=1$, since the kinetic energy term is tridiagonal.
@@
$|H(N,K)|=\begin{cases}\operatorname{d}(n,1)+n2^{n-1}&K=0\\ \operatorname{d}(n,K)+2^{n-1}\sum\limits_{k=1}^{K}\overline{n}_{k}&1\leq K\leq 2^{n-1}\\ 2^{n-1}(1+2^{n})&K>2^{n-1},\end{cases}$ (66)
as depicted in Fig. 4, where
$\operatorname{d}(n,K)=\sum_{m=0}^{K}\binom{n}{m}$, (67)
and $\overline{n}_{k}\coloneqq n-\lceil\operatorname{log}_{2}(k)\rceil$. The proof for this can be found in Lemma A.12. For the example of $N=4,K=2$, we see that the number of terms is $10$, confirmed by Eqs. (45) and (58). For $K>N/2$, we see that the number of Pauli terms is $\mathcal{O}(N^{2})$.
@@
$H_{N,K}$ Binary Gray
$K\leq N/2$ $K>N/2$ $K\leq N/2$ $K>N/2$
Qubits $\operatorname{log}_{2}(N)$ $\operatorname{log}_{2}(N)$
Pauli Terms $\operatorname{d}(n,K)+2^{n-1}\sum\limits_{k=1}^{K}\overline{n}_{k}$ $2^{n-1}(1+2^{n})$ $\operatorname{d}(n,K)+2^{n-1}\sum\limits_{k=1}^{K}\overline{n}_{k}$ $2^{n-1}(1+2^{n})$
QC Sets $1+\sum\limits_{k=1}^{K}2^{|b(\bar{k})|}[1-2^{-\overline{n}_{k}}]$ $\frac{1}{2}\left(1+3^{n}\right)$ $1+\sum\limits_{k=1}^{K}\overline{n}_{k}2^{|g(k-1)|}$ $\frac{1}{2}\left(1+3^{n}\right)$
DGC Sets $1+\sum\limits_{k=1}^{K}\overline{n}_{k}$ $2^{n}$ $1+\sum\limits_{k=1}^{K}\overline{n}_{k}$ $2^{n}$
2QG in Diag. Unitary $\frac{1}{2}\sum\limits_{k=1}^{K}\overline{n}_{k}\left[2|b(\overline{k})|-1-\overline{n}_{k}\right]$ $1+2^{n-1}(n-2)$ $\sum\limits_{k=1}^{K}\overline{n}_{k}|g_{k-1}|$ $1+2^{n-1}(n-2)$
Ansatz nLnL one-qubit gates + (n1)L(n-1)L two-qubit gates nLnL one-qubit gates + (n1)L(n-1)L two-qubit gates
Table 2: Number of qubit, Pauli terms, qubit-wise commuting sets and distance-grouped commuting sets for both the binary and Gray code, for a general Hamiltonian matrix of $2K+1$ bandwidth. Here we use $N=2^{n}$ and $\overline{n}_{k}\coloneqq n-\lceil\operatorname{log}_{2}(k)\rceil$. The quantity $\operatorname{d}(n,K)$ is defined in (67). For the ansatz, $L$ refers to the number of layers, which is chosen large enough beforehand. In the present study, the Hamiltonian for $K=0$ (diagonal potential) is tridiagonal because of the kinetic energy term, in which case the entries for $K=1$ should be used. Notably, for $K>N/2$, there is a saturation in all quantities, implying that larger diagonal width (better approximation) can be handled without an increase in the complexity of the problem. More details can be found in Sec. VI.
@@ V.2 Number of commuting sets
In addition, in this paper, we introduce a new strategy that is based on grouping Pauli strings in terms of the distance operators defined in Appendix A. We refer to this grouping as distance-grouped commutativity (DGC). While the precise structure is not relevant here, it leads to a more optimal set of Pauli operators as compared to the qubit-wise scheme. However, the diagonalizing unitary, while simple conceptually, is no longer a tensor-product of individual Pauli operators (see Appendix A for further details).
@@ V.2.4 Binary/Gray encoding: Distance-grouped commutativity
$=\begin{cases}0.5n(n-1)&K=0\\ 0.5\sum\limits_{k=1}^{K}\overline{n}_{k}\left[2|b(\overline{k})|-1-\overline{n}_{k}\right]&1\leq K\leq 2^{n-1}\\ 1+2^{n-1}(n-2)&K>2^{n-1},\end{cases}$ (81)
@@ VI Quantum simulations and discussions
[…]
Figure 9: Energy of the lowest $\frac{1}{2}^{+}$ state in (a) n+10C, (b) n+12C, (c) n+14C, (d) n+16C, and (e) n+18C, vs. the model-space size NN for $V_{\rm E}$ (13) (blue filled) and its $K=3$ approximation (purple open, for NN qubits; red open, for $n=\log_{2}N$ qubits), as compared to experiment (dashed line) [33, 2, 1]. Shaded area provides the error associated with the $K=3$ approximation (the small error band for n+10C is not visible on the plot).
@@ VI.1 Exponential potentials for neutron-Carbon dynamics
AA $E_{\rm expt}$ [MeV] $E_{\rm th}$ [MeV] $E_{\rm th}^{K=3}$ [MeV] $\frac{1}{\hbar\omega}V_{0}$ $c^{-\frac{1}{2}}$
n+10C $10$ 6.78-6.78 6.78-6.78 6.74-6.74 0.650-0.650 $5.43$
n+12C $12$ 1.86-1.86 1.86-1.86 1.74-1.74 0.283-0.283 $5.35$
n+14C $14$ 1.22-1.22 1.22-1.22 1.04-1.04 0.242-0.242 $5.0$
n+16C $16$ 0.52-0.52 0.52-0.52 0.23-0.23 0.175-0.175 $4.7$
n+18C $18$ 0.58-0.58 0.59-0.59 0.30-0.30 0.192-0.192 $4.6$
@@
For Gaussian-like exponential potentials, the truncation parameter $K$ needs to be odd, since for even $K$ values, the potential curves downward with increasing distance and leads to spurious bound states. Furthermore, the $K=3$ case provides a reasonable approximation, as illustrated by the $K=3$ error band in Fig. 9. This error is smaller for the lighter isotopes, where the $\frac{1}{2}^{+}$ state is more deeply bound (larger binding energy) compared to the one for n+16C and n+18C (see also $E_{\rm th}^{K=3}$ in Table 3). It is important to note that weakly bound states converge at a slower rate compared to states with larger binding energies. As shown in Fig. 9, the $N=8$ energy (or 3 qubits for the Gray encoding) closely agrees with the exact value for n+10C, but larger model spaces are needed for the other systems. The model-space size requirements become even larger for resonances. This suggests that the use of the Gray encoding that can reach large model spaces with fewer qubits becomes advantageous for weakly bound states and resonances. To illustrate the advantages of the Gray encoding compared to the one-hot encoding, we perform quantum simulations for n+14C using both encodings, which is detailed below.
[…]
@@
Figure 10: (a) Ab initio nonlocal n+4He potential as a function of the distance $r$ from the center-of-mass of the $\alpha$ particle for the $S_{\frac{1}{2}}$ partial wave, calculated in the SA-NCSM/GF with $\hbar\omega=16$ MeV and 15 HO shells, for zero projectile energy (see also Fig. 4 in Ref. [11] for other energies and partial waves). (b) & (c) Approximate potentials for different truncation parameters $K$, as compared to the ab initio deduced local potential labeled as “Local VV” (black, solid) for (b) $\hbar\omega=12$ MeV and (c) $\hbar\omega=16$ MeV. The corresponding exact lowest eigenvalue $E_{0}$ is also shown (gray, dotted).
@@
with the corresponding lowest eigenvalue of the original nonlocal potential $E_{0}=-20.84$ MeV. Hence, across the range of $\hbar\omega=12$-16 MeV, the ab initio potential yields the lowest $1/2^{+}$ orbit at energy $E_{0}=-19.8\pm 1.0$ MeV, which closely agrees with the corresponding 3He experimental energy of 20.58-20.58 MeV associated with a neutron removal from the 4He target.
@@
[…]